# The Butterfly Effect of Altering Prompts: How Small Changes and Jailbreaks Affect Large Language Model Performance

### v2

@@ Abstract
Large Language Models (LLMs) are regularly being used to label data across many domains and for myriad tasks. By simply asking the LLM for an answer, or “prompting,” practitioners are able to use LLMs to quickly get a response for an arbitrary task. This prompting is done through a series of decisions by the practitioner, from simple wording of the prompt, to requesting the output in a certain data format, to jailbreaking in the case of prompts that address more sensitive topics. In this work we ask: do variations in the way a prompt is constructed change the ultimate decision of the LLM? We answer this using a series of prompt variations across a variety of text classification tasks. We find that even the smallest of perturbations, such as adding a space at the end of a prompt, can cause the LLM to change its answer. Further, we find that requesting responses in XML and commonly-used jailbreaks can have cataclysmic effects on the data labeled by LLMs. Our code is available at https://github.com/Abel2Code/The_Butterfly_Effect_of_Prompts.
@@ 1 Introduction
In this work, we ask the question: How reliable are LLMs’ responses to variations in the prompts? We explore three types of variations in isolation. The first variation is to ask the LLM to give its response in a certain “output format.” Following common practice Li et al. (2023); Lee et al. (2023); Hada et al. (2023), [^1] Libraries exist to facilitate this, e.g., https://github.com/1rgs/jsonformer. we ask the LLM to format its output in frequently-used data formats such as a Python list or JSON. These are enumerated in Section 3.2.1. Second, we extend one of these formats–the Python list–and explore minor variations to the prompt. Fully enumerated in Section 3.2.2, these are small changes to the prompt such as adding a space, ending with “Thank you,” or promising the LLM a tip. [^2] These are only promised in the text. LLMs do not yet accept tips. The final type of variation we explore are “jailbreaks.” Practitioners wishing to label data concerning sensitive topics, like hate speech detection, often need to employ jailbreaks to bypass the LLM’s content filters. This practice has become so common that websites have emerged to catalog successf instances of this variation. [^3] E.g., https://www.jailbreakchat.com/ Listed in Section 3.2.3, we explore several commonly-used jailbreaks.
@@ 3.3 Experimental Setup
We conducted our experiments using OpenAI’s ChatGPT (gpt-3.5-turbo-1106). We opted for this model due to its widespread usage, public accessibility, and advanced generation capabilities. To ensure deterministic outputs, we set the temperature parameter to 0 which favors the selection of tokens with the highest probabilities at each step. It’s important to note that while this favors high-probability token selection at each step, it doesn’t guarantee the final sequence will have the highest overall probability. Nevertheless, this setting enables us to explore the model’s tendency to provide highly probable responses. Additionally, a temperature of 0 is often preferred in production settings due to its deterministic nature, which ensures consistency in generated outputs, and enables greater reproducibility. [^5] https://huyenchip.com/2023/04/11/llm-engineering.html We automatically parse model outputs, even attempting to parse incorrectly formatted results (e.g. JSON-like outputs that are technically invalid). These experiments were conducted from December 1st, 2023 to January 3rd, 2024.
@@ 4.1 Are predictions sensitive to prompt variations?
Yes! First, we analyze the impact of formatting specifications on predictions. In Figure 1, we demonstrate that by simply adding a specified output format, we observe a minimum of 10% of predictions change. Notably, even just utilizing the ChatGPT’s JSON Checkbox feature via the ChatGPT API results in even more prediction changes compared to simply using the JSON specification.
@@
While the impact of our perturbations is smaller than changing the entire output format, a significant number of predictions still undergo change. Intriguingly, even introducing a simple space at the prompt’s beginning or end leads to over 500 prediction changes. Similar effects are observed when adding common greetings or ending with “Thank you”. Among the perturbations, rephrasing as a statement exhibits the most substantial impact, albeit affecting over 900 predictions.
We observe that using jailbreaks on these tasks leads to a much larger proportion of changes. Notably, AIM and Dev Mode V2 yield invalid responses in around 90% of predictions, primarily due to the model’s standard response of “I’m sorry, I cannot comply with that request.” Despite the innocuous nature of the questions used with the jailbreaks, we suspect that ChatGPT’s fine-tuning specifically avoids responding to these jailbreaks.
While we find much more success in receiving valid responses with Refusal Suppression and Evil Confidant, their usage results in over 2500 prediction changes just by including these jailbreaks. Evil Confidant, expectedly, prompts a significant shift, given its directive for the model to provide “unhinged” answers. We expected less shift when using Refusal Suppression, yet it also yielded a substantial deviation in predictions.
Figure 1 aggregates the changes across all 11 tasks. For number of prediction changes on a per-task level, see Appendix B.

### v3

@@ Abstract
Large Language Models (LLMs) are regularly being used to label data across many domains and for myriad tasks. By simply asking the LLM for an answer, or “prompting,” practitioners are able to use LLMs to quickly get a response for an arbitrary task. This prompting is done through a series of decisions by the practitioner, from simple wording of the prompt, to requesting the output in a certain data format, to jailbreaking in the case of prompts that address more sensitive topics. In this work we ask: do variations in the way a prompt is constructed change the ultimate decision of the LLM? We answer this using a series of prompt variations across a variety of text classification tasks. We find that even the smallest of perturbations, such as adding a space at the end of a prompt, can cause the LLM to change its answer. Further, we find that requesting responses in XML and commonly-used jailbreaks can have cataclysmic effects on the data labeled by LLMs.
@@ 1 Introduction
In this work, we ask the question: How reliable are LLMs’ responses to variations in the prompts? We explore three types of variations in isolation. The first variation is to ask the LLM to give its response in a certain “output format.” Following common practice Li et al. (2023); Lee et al. (2023); Hada et al. (2023), [^1] Libraries exist to facilitate this, e.g., https://github.com/1rgs/jsonformer. we ask the LLM to format its output in frequently-used data formats such as a Python list or JSON. These are enumerated in Section 2.2.1. Second, we extend one of these formats–the Python list–and explore minor variations to the prompt. Fully enumerated in Section 2.2.2, these are small changes to the prompt such as adding a space, ending with “Thank you,” or promising the LLM a tip. [^2] These are only promised in the text. LLMs do not yet accept tips. The final type of variation we explore are “jailbreaks.” Practitioners wishing to label data concerning sensitive topics, like hate speech detection, often need to employ jailbreaks to bypass the LLM’s content filters. This practice has become so common that websites have emerged to catalog successful instances of this variation. [^3] E.g., https://www.jailbreakchat.com/ Listed in Section 2.2.3, we explore several commonly-used jailbreaks.
@@ 2.3 Experimental Setup
We conducted our experiments using OpenAI’s ChatGPT (gpt-3.5-turbo-1106) and all variations of Llama 2 (7B, 13B, and 70B). We opted for these models due to their widespread usage, public accessibility, and advanced generation capabilities.
To ensure deterministic outputs, we set the temperature parameter to 0 which favors the selection of tokens with the highest probabilities at each step. It’s important to note that while this favors high-probability token selection at each step, it doesn’t guarantee the final sequence will have the highest overall probability. Nevertheless, this setting enables us to explore the model’s tendency to provide highly probable responses. Additionally, a temperature of 0 is often preferred in production settings due to its deterministic nature, which ensures consistency in generated outputs, and enables greater reproducibility. [^5] https://huyenchip.com/2023/04/11/llm-engineering.html
We automatically parse model outputs, even attempting to parse incorrectly formatted results (e.g. JSON-like outputs that are technically invalid). These experiments were conducted from December 1st, 2023 to January 3rd, 2024.
@@ 3.1 Are predictions sensitive to prompt variations?
Yes! First, we analyze the impact of formatting specifications on predictions. In Figure 1, we demonstrate that by simply adding a specified output format, we observe a minimum of 10% of predictions change. Notably, even just utilizing ChatGPT’s JSON Checkbox feature via the ChatGPT API results in even more prediction changes compared to simply using the JSON specification.
@@
While the impact of our perturbations is smaller than changing the entire output format, a significant number of predictions still undergo change. Intriguingly, even introducing a simple space at the prompt’s beginning or end leads to over 500 prediction changes in ChatGPT. Llama’s implementation automatically strips input, thus the tokenized input is the same as the baseline. We observed that even common greetings or ending with “Thank you” changed a large amount of predictions. Among the perturbations, rephrasing as a statement typically exhibited the most substantial impact.
We observe an interesting trend with regard to model size. As the number of parameters increases, the models seemingly become more robust to these variations. This behavior is unsurprising. When the model has fewer parameters, we would expect more reliance on spurious correlations, like our variations, having more impact on the final output.
We observe that using jailbreaks on these tasks leads to a much larger proportion of changes overall. Notably, AIM and Dev Mode V2 yield invalid responses in around 90% of predictions for ChatGPT, primarily due to the model’s standard response of “I’m sorry, I cannot comply with that request.” Despite the innocuous nature of the questions used with the jailbreaks, we suspect that ChatGPT’s fine-tuning specifically avoids responding to these jailbreaks. Surprisingly, Llama saw opposite behavior with the number of invalid responses decreasing as the parameter size increased.
We saw the opposite behavior for Refusal Suppression and Evil Confidant, where invalid response frequency increased with parameter size in Llama, yet ChatGPT saw few invalid responses. The mere inclusion of these jailbreaks results in over 2500 prediction changes (out of 11000) for ChatGPT alone, the largest amount of changes in ChatGPT compared to any other variation. Evil Confidant, expectedly, prompts a significant shift, given its directive for the model to provide “unhinged” answers. We expected less shift when using Refusal Suppression, yet it also yielded a substantial deviation in predictions.
Figure 1 aggregates the changes across all $1$ tasks. The number of prediction changes on a per-task level is reported in Appendix B.
@@ 3.2 Do prompt variations affect accuracy?
Llama-7B
Llama-13B
Llama-70B
ChatGPT
Python List Format 41.8% 57.7% 65.0% 78.6%
JSON Format 46.1% 56.4% 68.8% 78.5%
ChatGPT’s JSON Checkbox N/A N/A N/A 73.2%
XML Format 43.7% 54.7% 56.2% 74.4%
CSV Format 42.1% 57.4% 63.9% 73.2%
YAML Format 43.5% 57.4% 61.4% 76.7%
No Specified Format 42.2% 53.7% 65.2% 79.6%
Start with Space N/A N/A N/A 78.5%
End with Space N/A N/A N/A 78.4%
Start with "Hello." 42.9% 54.9% 63.3% 78.0%
Start with "Hello!" 43.8% 56.1% 64.2% 78.0%
Start with "Howdy!" 39.7% 54.6% 62.6% 78.0%
End with "Thank you" 43.1% 56.5% 64.4% 78.0%
Rephrase as Statement 49.4% 54.4% 64.3% 78.3%
Won’t Tip 35.3% 55.2% 63.1% 78.0%
Tip \$1 52.0% 57.9% 62.1% 78.2%
Tip \$10 52.6% 56.1% 61.0% 78.3%
Tip \$100 50.6% 54.0% 59.0% 78.2%
Tip \$1000 47.8% 52.0% 56.9% 78.1%
AIM 19.3% 30.1% 55.0% 6.3%
Evil Confidant 29.0% 20.5% 18.0% 60.4%
Refusal Suppression 42.6% 55.0% 56.5% 67.1%
Dev Mode v2 26.4% 46.3% 45.0% 4.1%
Aggregate Output Formats 48.5% 59.5% 69.3% 79.9%
Aggregate Perturbations 45.4% 57.1% 65.1% 78.7%
Aggregate Jailbreaks 35.1% 38.5% 56.3% 51.3%
Aggregate Tipping 51.6% 55.8% 60.9% 78.8%
Table 1: Overall accuracy of each prompt variation across all tasks.
Yes! Table 1 shows the accuracy of each prompt variation across all 4 models. There is no task that objectively outperforms the others across all tasks or models, although we generally observe success using the Python List, No Specified Format, or JSON specification. No Specified Format leads to the overall most accurate results on ChatGPT, beating the next best variation by a whole percentage point. Llama, on the other hand, performs best with the JSON formatting constraint on Llama-7B and Llama 70B,, however this does slightly worse than other formats for Llama-13B.
Formatting in YAML, XML, or CSV do worse compared to No Specified Format for our largest models, Llama-70B and ChatGPT. Llama-7B and 13B interestingly see an increase in performance for these variations. These improvements or degradations are not necessarily consistent across tasks. For example, CSV is the worst performing style variation (tied with ChatGPT’s JSON Checkbox)) yet it achieves the highest accuracy among all variations for the IMDBSentiment task, albeit by only a marginal percentage point. This emphasizes the absence of a definitive “best” or “worst” output format for usage.
When it comes to influencing the model by specifying a tip versus specifying we will not tip, we found that tipping \$1, \$10, or \$100 to Llama-7B significantly improves the performance, outperforming every other variation we tested. This performance increase is not seen in larger models tested. We saw minimal differences in performance in ChatGPT when tipping versus not. This suggests that larger models are more robust to spurious tokens in classification tasks. Contrary to expectations, tipping extravagant amount to any model, specifically \$1000, led to degradation in accuracy compared to tipping less.
Furthermore, our experimentation revealed a significant performance drop when using certain jailbreaks. AIM and Dev Mode v2 unsurprisingly exhibit very low accuracy for ChatGPT, primarily due to a majority of their responses being invalid. Given that Llama-2 saw less invalid responses as the model size increased, AIM’s performance improved with model size, although Llama-13B and Llama-70B saw similar performance for Dev Mode V2. Evil Confidant, with its prompt guiding it toward “unhinged” responses, also yields low accuracy overall. Surprisingly, the Refusal Suppression resulted in an over 9% loss in accuracy (compared to Python List) for both Llama-70B and ChatGPT, highlighting the inherent instability even in seemingly innocuous jailbreaks. We do, however, see only a 2% decrease in accuracy for Llama-13B and a slight increase for Llama-7B. This underscores the unpredictability associated with jailbreak usage.
We additionally explored the effects of majority voting. Self-consistency (Wang et al., 2023) is a technique that prompts a model multiple times, with a non-zero temperature and the same prompt, and uses the most common prediction as a final answer. We aggregate our predictions across prompt variations, rather than resampling with a larger temperature. One benefit of this approach is that it is able to generate predictions despite some of the variations returning invalid responses. We find that this approach provides clear benefits to the overall accuracy, with Aggregate Output Formats achieving the highest overall accuracy across all models, except Llama 7B, where it was beaten only by the tipping strategy.
3.3 How similar are the predictions from each prompt variation?
(a) Llama 7B
(b) Llama 13B
(c) Llama 70B
(d) ChatGPT
Figure 3: MDS representation of model predictions on prompt variations. Each prompt variation is encoded as a vector, with each dimension representing its corresponding response across all tasks. In this vector, ’1’ signifies correct predictions, ’-1’ indicates incorrect predictions, and ’0’ denotes invalid predictions.
First, we observe an interesting relationship in ChatGPT between Python List specification and the No Specified Format. These two vectors are placed close together in the MDS representation. We note again that these two formats also achieved the highest overall accuracy for ChatGPT. This relationship does not stay true for our Llama models. Adjacent to these points in ChatGPT were simple perturbations, which were formatted as Python lists, such as initial greetings or the addition of a space. This clustering around the Python List variation may be attributed to these prompts having only a few token differences while preserving the overall semantics, although this relationship was more variable across Llama models.
Contrary to expectations, all tipping variations clustered together across all models, with even the Won’t Tip variation being included in this cluster for ChatGPT. Surprisingly, increasing the tip amount exhibited a linear relationship with distances from the Won’t Tip variation in ChatGPT.
Rephrase as Statement stood out as an outlier across all models, situated far from the main clusters. The substantial impact of rephrasing was expected, given the increased token changes compared to other prompts. End with "Thank you" additionally stood as an outliar for ChatGPT. It is surprising that simply thanking the model can lead to such a considerable difference, while adding a greeting or space token leads to a minimal change.
Lastly, the jailbreak variations displayed a wider spread. These variations would often lead to invalid responses, aligning with their broader distribution. Surprisingly, Refusal Suppression fell on the outskirts of the primary cluster in ChatGPT’s representation, possibly due to the extensive token addition through the jailbreak. Despite requiring fewer tokens, the Evil Confidant variation notably diverged from the cluster main clusters as well, which we attribute to its directive to produce “unhinged” responses.
Category ChatGPT Llama-7B Llama-13B Llama-70B
All -0.2334 (p = 0.00) -0.3674 (p = 0.00) -0.2686 (p = 0.00) -0.1509 (p = 0.00)
Styles -0.0669 (p = 0.03) -0.2328 (p = 0.00) -0.1676 (p = 0.00) -0.0786 (p = 0.01)
Perturbations 0.1209 (p = 0.00) -0.2307 (p = 0.00) -0.0437 (p = 0.17) 0.1541 (p = 0.00)
Tipping 0.1241 (p = 0.00) -0.1614 (p = 0.00) -0.1537 (p = 0.00) 0.0909 (p = 0.00)
Jailbreaks -0.3779 (p = 0.00) -0.3578 (p = 0.00) -0.3536 (p = 0.00) -0.4047 (p = 0.00)
3.4 Do variations correlate to annotator disagreement?
Not really! Leveraging the Jigsaw Toxicity task, which we specifically sampled only to include samples with 100 or more annotations, we hypothesized that more confusing samples would lead to more annotator disagreement and more variation in our model’s predictions. To aid our analysis, we calculate the entropy of annotator predictions and the entropy of our predictions per sample.
Table 2 lists the Pearson correlations between the Jigsaw Toxicity predictions, across each category of prompt variations. We identify some weak correlations with annotator disagreement. However, the strongest correlations are negative, meaning that the least confusing instances (i.e., lowest entropy) and the most likely to change. This indicates that the confusion of the instance provides some explanatory power for why the prediction changes, but there are other factors at play.
4 Related Work
Sclar et al. (2023) investigated the sensitivity of LLMs to arbitrary prompt formatting choices in few-shot settings, like capitalization or changes in prompt formatting such as varying capitalization or word choice in formatting context of a prompt (i.e. “Passage: ” vs “Context: ”). They identified performance differences across models. Our work focuses on a broader range of variations, which contain semantic meaning that should not effect the expected answer. Additionally, our work differentiates by investigating the effects of output formats in predictions, a commonly used prompting strategy for LLM evaluation.
Bsharat et al. (2023) examined the effectiveness of various prompting “principles” in ChatGPT and Llama, with the goal of providing practitioners with suggested prompt strategies. Among their recommendations were to avoid phrases like "please" and “‘thank you” and to add “‘I’m going to tip \$xxx for a better solution!”. They measure the effectiveness of these principles by having human evaluators judge the quality and correctness of LLM responses. They find significant improvements across all models when using the principles highlighted above. Our analysis, which evaluates perturbations through classification tasks with ground truth labels, instead finds the effectiveness of the tipping principle and removing thank you to be much smaller than the results showcased in their paper, with the exception of tipping having a large effect on Llama-7B.
5 Conclusion
In this paper, we investigate how simple and commonly-used prompt variations can affect an LLM’s predictions. We demonstrate that even minor prompt variations can change a considerable proportion of predictions. That said, despite some fraction of labels changing, most perturbations yield similar accuracy. We find that jailbreaks lead to considerable performance losses. The AIM and Dev Mode v2 jailbreaks led to refusal rates around 90% for ChatGPT. Additionally, while both Evil Confidant and Refusal Suppression had a refusal rate of less than 3%, their inclusion led to a loss of over 10 percentage points compared to our baseline. Finally, we observe a performance hit when using specific output format specifications, a commonly used approach for classification evaluation.
Next, we analyze the patterns of these changes. First, we embed the prompt variations based on their subsequent responses using MDS, and find that perturbation outputs tend to more closely resemble our baseline than formatting changes, and that both have higher fidelity than jailbreaks. Next, we study the correlation between annotator disagreement and an instance’s propensity to change. We find a slight negative correlation between annotator disagreement and the likelihood to change.
The directions for future work are abundant. A major next step would be to generate LLMs that are resilient to these changes, offering consistent answers across formatting changes, perturbations, and jailbreaks. Towards that goal, future work includes seeking a firmer understanding of why responses change under minor changes to the prompt, and better anticipating an LLMs change in its response to a particular instance.
6 Limitations
Our study delves into the impact of minor variations in prompts on the predictions and overall performance of large language models. While we explore a wide array of prompt variations, it’s crucial to note that even within our prompt variations, we followed some consistent wordings or formatting styles (such as delimiter choice). These choices can have discernible effects on the models’ performance or predictions.
Moreover, we observed that the relative performance of prompt variations could differ significantly across various classification tasks. Our analysis primarily focuses on classification tasks; however, future research endeavors could extend this investigation to explore prompt sensitivity in scenarios involving open-ended questions or short-answer tasks.
Finally, our examination is constrained to two specific model variations, namely ChatGPT and Llama. It is imperative to conduct further investigations to comprehend how different models, architectures, training data, and other factors may influence the sensitivity of models to prompt variations. Such investigations would offer a more comprehensive understanding of the broader implications of prompt engineering on model behavior and performance.
Appendix A Full Prompts
A.1 Tasks
Each task and a corresponding example prompt is shown in Table 3.
Task Prompt
Table 3: Examples of each task’s prompt.
A.2 Variations
Table 4: Examples of each variation’s prompt.
Variation Example
Appendix B Extended Results
For completeness, we include more granular results of the experiments presented in our paper. Table  9, 10, 11, 12 present the number of style predictions that change from No Specified Format for each individual dataset. Table  13, 14, 15, 16 present the number of style predictions that change from No Specified Format for each individual dataset. Table 5, 6, 7, and 8 presents the accuracy on a per dataset level. In our paper, we discussed how many overall predictions change when a prompt variation is used.
@@ Appendix B Extended Results
Aggregate Output Formats 83% 84% 82% 93% 47% 92% 61% 87% 88% 83% 78% 80%
Aggregate Perturbations 78% 83% 79% 92% 47% 92% 61% 90% 85% 82% 77% 79%
Aggregate Jailbreaks 55% 50% 48% 57% 41% 78% 28% 66% 20% 74% 48% 51%
Aggregate Tipping 77% 83% 80% 94% 49% 92% 57% 91% 85% 83% 77% 79%